# The Generalization Ridge: Information Flow in Natural Language Generation

### v1

@@ 5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
Figure 5: Middle transformer blocks are key to encoding generalizable task-relevant information (GPT-2 Small, Synthetic).
To understand how information accumulates across the network, we compute Incremental Information Gain ($I(\Delta Z_{\ell};Y)$)—the mutual information between each residual transition and the target label embedding. Figure 5 shows the layer-wise gains, revealing that intermediate layers yield the highest information increases. This concentration of information gain highlights their central role in encoding task-relevant features essential for generalization. A case study decoding $\Delta z$ via the LM head is provided in Appendix F, highlighting token-level shifts across layers. The Incremental Information Gain analysis reveals a clear pattern: middle transformer blocks are key to encoding generalizable task-relevant information, thereby forming a generalization ridge. Conversely, later layers contribute little additional predictive signal, and in some cases, actively reduce alignment with the target embeddings. This diminishing contribution in later layers may suggest a shift from general reasoning to memorization. This trend underscores a fundamental trade-off in transformer training dynamics.
@@ 6 Residual Scaling Dynamics via Learnable $\beta$ Coefficients
Post-peak residual blocks (i.e., layers beyond the generalization ridge) may encode memorized signals. Suppressing these layers’ contribution via learned residual scaling improves out-of-distribution (OOD) generalization, while amplifying them degrades it.
We introduce a residual scaling mechanism with learnable scalar coefficient parameters (Algorithm 1), inspired by prior work on adaptive residual modulation liu2019self ; menghani2024laurel . Transformer architectures inherently employ residual connections to iteratively refine representations. To isolate and quantify the contribution of each transformer blocks, we scale these residual connections with layer-specific scaling factors $\beta_{\ell}\in\mathbb{R}_{\geq 0}$, with definition below. Each $\beta_{\ell}$ controls the strength of the residual contribution from layer $\ell$, enabling the model to adaptively emphasize or suppress specific blocks:
@@
We freeze model weights and optimize only the residual scaling coefficient parameters $\beta_{\ell}$, which are initialized to 1. These scalars are trained on the (a) in-distribution (ID) split and (b) out-of-distribution (OOD) split. Since no other parameters are updated, the learned $\beta_{\ell}$ values serve as a direct diagnostic of the extent to which each layer should be amplified or attenuated to suit the data regime.
As shown in Table 2, optimizing the residual scaling coefficient parameters on in-distribution results in $\beta_{12}>1$, indicating that in-distribution performance benefits from amplifying the contribution of the final transformer layers—suggestive of their role in memorizing features specific to the training distribution. In contrast, when trained on out-of-distribution (OOD) data, the final layer’s contribution is downweighted ($\beta_{12}<1$), implying that suppressing memorization-heavy components improves generalization. This pattern holds consistently across architectures such as Qwen and LLaMA (Figure 6) and corresponds to improved OOD performance, as shown in Table 3. Together, these findings provide empirical support for the functional specialization of transformer layers: intermediate layers tend to encode generalizable representations, whereas deeper layers increasingly capture memorized signals.
Table 2: $I(Z;Y)$ and residual scaling coefficients $\beta_{\ell}$ for GPT-2 Small. Optimizing on in-distribution data yields $\beta_{12}>1$, while OOD training suppresses the final layer’s contribution with $\beta_{12}<1$.
Layer 1 Layer 2 Layer 3 Layer 4 Layer 5 Layer 6 Layer 7 Layer 8 Layer 9 Layer 10 Layer 11 Layer 12
$I(Z;Y)$ (CLUTRR) 0.0526 0.0474 0.0407 0.0372 0.0319 0.0337 0.0837 0.2688 0.3430 0.3901 0.4514 0.0851
$\beta_{\ell}$ (CLUTRR ID) 0.9998 0.9999 1.0002 1.0002 1.0003 1.0004 0.9993 0.9997 0.9998 0.9993 0.9999 1.0003
$\beta_{\ell}$ (CLUTRR OOD) 1.0005 1.0005 1.0000 1.0000 0.9999 0.9997 1.0002 1.0003 1.0002 0.9995 0.9996 0.9995
$I(Z;Y)$ (Synthetic) 0.0508 0.0513 0.0619 0.0822 0.0894 0.0989 0.1292 0.2431 0.2810 0.3209 0.3209 0.1402
$\beta_{\ell}$ (Synthetic ID) 0.9993 1.0007 0.9995 1.0005 1.0000 0.9999 1.0000 0.9999 0.9995 1.0018 1.0003 1.0001
$\beta_{\ell}$ (Synthetic OOD) 1.0015 1.0003 0.9974 1.0006 1.0010 0.9980 0.9987 0.9974 0.9985 0.9982 0.9981 0.9956
Table 3: Loss and accuracy (%) before and after residual scaling on out-of-distribution (OOD) split. Performance improves following the optimization of residual scaling coefficient parameters.
Dataset Clutrr Synthetic
[…]
GPT-2 Small 6.9655 21.24 6.9636 21.24 2.6533 70.53 2.6485 70.54
Qwen-2.5-0.5B 4.2494 31.01 4.2465 31.01 1.8132 74.31 1.7217 74.55
LLaMA-3.1-8B 1.3209 57.23 1.3193 57.13 1.3927 87.39 1.3155 87.83
Figure 6: Residual scaling coefficients $\beta_{\ell}$ for Qwen-2.5-0.5B and LLaMA-3.1-8B. A similar pattern consistently holds across architectures, with $\beta_{\text{final layer}}$ decreasing under OOD training.
7 Case Study: Attention Dynamics Across Layers
@@ 8 Conclusion
We present InfoRidge, an information-theoretic framework, to analyze how information evolves across layers in transformer language models. By estimating both predictive information and incremental information gain, we systematically characterize the dynamics of task-relevant signal propagation throughout the network. Our findings reveal a consistent generalization ridge emerging in intermediate layers, where mutual information between the hidden representation and the target label reaches its peak. This phenomenon reflects a fundamental trade-off between generalization and memorization as information flows deeper into the model. Residual scaling experiments further corroborate this interpretation, demonstrating the functional specialization of different layers—where intermediate blocks play a key role in supporting generalization, while deeper layers increasingly focus on memorization. Attention flow provides additional insight into where the model concentrates information. Taken together, these results offer a unified perspective on information flow in transformers for natural language generation and point to opportunities for enhancing model architecture, improving interpretability, and increasing robustness to distributional shifts.
9 Limitation
Our analysis focuses on a select set of datasets and model architectures, which may limit the breadth of empirical coverage. The mutual information estimates rely on kernel-based approximations that may be sensitive to representation geometry. Additionally, the residual scaling approach assumes layer-wise independence and does not account for potential inter-layer dependencies during adaptation. Future work could address these limitations by extending the analysis to broader model families and tasks and exploring joint optimization strategies.
@@ Appendix A Mathematical Details for Matrix-Based Information Estimation and Theoretical Foundations
We employ the matrix-based Rényi entropy  [11] to estimate mutual information between model representations and target labels. This method captured sample similarity structure via kernel Gram matrices.
@@ A.1 Matrix-Based Entropy Estimation
Let $U=\{u_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$ denote $\ell_{2}$-normalized representations obtained from a specific transformer layer, a residual update, or the embedding of the target label. A Gaussian kernel Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is constructed as:
$(G_{U})_{ij}=\exp\left(-\frac{\|u_{i}-u_{j}\|^{2}}{2\sigma^{2}}\right),$
with bandwidth $\sigma=1$. The matrix is then trace-normalized to ensure $\text{tr}(G_{U})=1$.
@@ A.2 Mutual Information Estimation
To estimate the mutual information between two random variables $U$ and $V$, we compute their Gram matrices $G_{U}$ and $G_{V}$, and form the joint similarity matrix via element-wise (Hadamard) product:
$G_{UV}=G_{U}\circ G_{V}.$
After trace-normalization, mutual information is estimated by:
$I(U;V)=H(U)+H(V)-H(U,V),$
where $H(U,V)=-\text{tr}(G_{UV}\log G_{UV})$. The eigenvalue spectrum of $G_{UV}$ governs the joint entropy term; its shape reflects how much of the structure in $U$ and $V$ aligns. A more concentrated spectrum in $G_{UV}$ relative to $G_{U}$ and $G_{V}$ implies stronger dependence and thus higher mutual information.
A.3 Min Wasserstein Generalization Bound [20]
We restate the generalization bound proposed by He et al. [20], which characterizes generalization in terms of the Wasserstein distance between internal representations.
Suppose that the loss function ~$\tilde{\ell}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is $\rho_{0}$-Lipschitz, and the activation function $\phi_{\ell}:\mathbb{R}\to\mathbb{R}$ is $\rho_{\ell}$-Lipschitz for each $\ell=1,\ldots,L$. Then:
$\small\mathrm{gen}(P_{W|\mathcal{D}_{n}}\!,\!P_{X,Y})\!\leq\!\min_{\ell}\frac{\rho_{0}}{n}\sum_{i=1}^{n}\!\mathbb{E}_{W}\!\left[\!\left(1\!\vee\!\prod_{j=\ell+1}^{L}\!\rho_{j}\|W_{j}\|_{\mathrm{op}}\!\right)\!\cdot\!\mathcal{W}_{1}\!\left(\!P_{T_{\ell},i,Y_{i}|W}(\cdot|W),\!\ P_{T_{\ell},Y|W}(\cdot|W)\!\right)\!\right]$ (3)
This result shows that generalization can be tightly controlled by the Wasserstein distance [21] between internal representations at a specific layer—referred to as the generalization funnel layer.
@@ A.3 Min Wasserstein Generalization Bound [20]
The Min Wasserstein Generalization Bound [20] provides a theoretical foundation for our study by characterizing generalization in terms of distributional alignment at an intermediate layer. It motivates our analysis of information flow by suggesting that information is peaked at a specific layer—the generalization funnel. Our InfoRidge framework builds on this insight by quantifying predictive information across layers, and reveals that a specific intermediate layer exhibits peak mutual information and correlates with better generalization performance.
@@ B.2 CLUTRR
CLUTRR (Compositional Language Understanding and Text-based Relational Reasoning) [25] is a diagnostic benchmark for evaluating relational inference in language models. Each example contains a story describing family relations, and the task is to infer the missing relationship between two entities. The distribution shift stems from clause lengths that are absent in the training set but present during evaluation. We use the task split “gen_train23_test2to10”, where the model is trained on clause lengths 2 and 3 and evaluated on lengths 2 through 10.
@@ B.3 ECQA
ECQA (Explanations for CommonsenseQA) [26] is a commonsense multiple-choice question-answering dataset, where each question is accompanied by 5 answer options.
@@ B.4 Synthetic Arithmetic Dataset
[…]
@@ C.1 Prompt Construction
All tasks are cast into a next-token generation format. The model receives a prompt (context) and must generate the correct continuation token. Below are construction strategies and examples for each dataset:
@@ E. Drive
C.2 Information Estimation
To estimate mutual information, we subsample between 50 and 200 test examples, with the exact number determined by the model and task, in order to avoid excessive memory usage from large Gram matrices. Full results and statistical significance are reported in Appendix E.
C.3 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
@@
  • $\beta$ values are initialized to 1 and updated using gradient descent without altering model architecture.
@@ Appendix D Compute and Licensing Details
All experiments were conducted on an NVIDIA RTX A6000 GPU with 48GB of memory.
@@
The GPT-2 Small [22] and GPT-2 Medium [22] models are released under the MIT License and are available via Hugging Face Transformers. The Qwen-2.5 0.5B [23] model is provided by Alibaba Group under the Apache 2.0 License. The LLaMA-3.1 8B [24] model is made available by Meta under a non-commercial research license and accessed via Hugging Face.
@@
CLUTRR [25] is released under a CC BY 4.0 license as part of EMNLP 2019. The ECQA [26] dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021. The Synthetic Arithmetic dataset is custom-designed by the authors and does not rely on any external or licensed data sources.
@@ Appendix E Full Quantitative Results with Confidence Estimates
To provide a comprehensive view of model behavior across layers, we report the full experimental results for all models and datasets evaluated in this study. This includes predictive information trends ($I(Z;Y)$), incremental information gain ($I(\Delta Z;Y)$) and residual scaling effects ($\beta_{\ell}$), together with statistical significance reporting.
@@ E.1 Predictive Information
To estimate mutual information between hidden representations and target labels, we compute matrix-based mutual information following the methodology outlined in Appendix A. Due to the high memory cost of computing $N\times N$ Gram matrices, we subsample between 50 and 100 test examples per experiment and report results with statistical significance based on multiple random seeds.
@@
Figure 8: Predictive information $I(Z;Y)$ across different models and datasets exhibits an information peak, indicating a generalization ridge. In cases where the task is too simple relative to model capacity—such as the synthetic arithmetic task with LLaMA—this trend reflects an overfitting regime. Lighter line colors represent later training epochs. Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.

### v2

@@ 5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
To understand how information accumulates across the network, we compute Incremental Information Gain ($I(\Delta Z_{\ell};Y)$)—the mutual information between each residual transition and the target label embedding. As shown in Figure 5, the resulting layer-wise gains reveal that intermediate layers yield the highest information increases. This concentration of information gain further underscores their central role in encoding those task-relevant features that are essential for supporting generalization. For additional illustration, a detailed case study that decodes $\Delta z$ via the LM head is provided in Appendix G, highlighting fine-grained token-level shifts and variations observed across layers.
Figure 5: Middle blocks are key to encoding generalizable information (GPT-2-S, Synthetic).
The Incremental Information Gain analysis reveals a clear pattern: middle transformer blocks are key to encoding generalizable task-relevant information, thereby forming a generalization ridge. Conversely, later layers contribute little additional predictive signal, and in some cases, actively reduce alignment with the target embeddings. This diminishing contribution in later layers may suggest a shift from general reasoning to memorization of training-specific patterns. This trend underscores a fundamental trade-off in transformer training dynamics.
@@ 6 Residual Scaling Dynamics via Learnable $\beta$ Coefficients
Post-peak residual blocks (i.e., layers beyond the generalization ridge) may encode memorized signals. Suppressing these layers’ contribution via learned residual scaling improves out-of-distribution generalization, while amplifying them degrades it.
We introduce a residual scaling mechanism with learnable scalar coefficient parameters (Algorithm 1), inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024). Transformer architectures inherently employ residual connections to iteratively refine representations. To isolate and quantify the contribution of each transformer block, we scale these residual connections with layer-specific scaling factors $\beta_{\ell}\in\mathbb{R}_{\geq 0}$, with definition below. Each $\beta_{\ell}$ controls the strength of the residual contribution from layer $\ell$, enabling the model to adaptively emphasize or suppress specific blocks:
@@
We freeze model weights and optimize only the residual scaling coefficient parameters $\beta_{\ell}$, which are initialized to 1. These scalars are trained separately on the (a) in-distribution (ID) split and (b) out-of-distribution (OOD) split. Since no other parameters are updated, the learned $\beta_{\ell}$ serve as a direct diagnostic of the extent to which each layer should be amplified or attenuated to suit the data regime, revealing which layers remain stable across regimes and which adapt strongly to distributional shifts.
(a) GPT-2 Small on Synthetic Arithmetic
(b) GPT-2 Small on CLUTRR
Figure 6: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. ID training emphasizes later layers, while OOD training has lower final layer weight. Curve shows the mean across 5 independent run, and the shaded region denotes 1-$\sigma$ error bar.
As shown in Figure 6, optimizing the residual scaling coefficient parameters on in-distribution data assigns relatively higher residual weights to the final transformer layers, indicating that in-distribution performance benefits from increased reliance on late-layer residual contributions. This observation suggests that these deeper layers capture refinements that are closely tied to the training distribution. In contrast, under out-of-distribution (OOD) settings, the learned coefficients systematically have relative lower contribution of the final layers compared to the ID settings, revealing that improved generalization is associated with downweighting late-stage residual signals. More results are in Appendix E. This pattern supports the interpretation that late-layer residuals become increasingly specialized in memorized patterns tied to the training distribution. Together, these findings provide empirical support for the generalization ridge hypothesis, demonstrating that information flow in transformers reflects a trade-off between generalizable signals and memorized, distribution-specific features.
@@ 7 Conclusion
We introduce InfoRidge, an information-theoretic framework designed to trace and quantify how information evolves across layers in transformer-based language models for natural language generation. By estimating both predictive information and incremental information gain, we systematically characterize the layerwise dynamics of information flow, offering a principled view of how signals are refined, amplified, or diminished as they propagate through the network. Our findings reveal a consistent generalization ridge emerging in intermediate layers, where mutual information between the hidden representation and the target label reaches its peak before gradually declining. This phenomenon reflects a fundamental trade-off between generalization and memorization as information flows deeper into the model. Residual scaling experiments further corroborate this interpretation, demonstrating the functional specialization of different layers. Taken together, these findings position InfoRidge as a comprehensive framework for diagnosing how language models internally manage information during natural language generation, while also revealing the structural mechanisms that govern the balance between generalization and memorization in transformer architectures.
Ethics Statement.
This work adheres to the Code of Ethics. Our experiments use only open-source models and publicly available datasets under their respective open licenses, with no involvement of human subjects or sensitive data. We identify no foreseeable ethical risks.
Reproducibility Statement.
We ensure reproducibility by providing experimental and implementation details in Section 4 and Appendices BC. Full results with statistical significance are in Appendix E, and anonymous source code is included as supplementary material.
@@ Appendix A Mathematical Details for Matrix-Based Information Estimation and Theoretical Foundations
We employ the matrix-based Rényi entropy  Giraldo et al. (2014) to estimate mutual information between representations and labels, leveraging kernel Gram matrices to capture sample similarity.
@@ A.1 Matrix-Based Entropy Estimation
Let $U=\{u_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$ denote $\ell_{2}$-normalized representations obtained from a specific transformer layer, a residual update, or the embedding of the target label. A Gaussian kernel Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is constructed as: $(G_{U})_{ij}=\exp\left(-\frac{\|u_{i}-u_{j}\|^{2}}{2\sigma^{2}}\right),$ with bandwidth $\sigma=1$. The matrix is then trace-normalized to ensure $\text{tr}(G_{U})=1$.
@@ A.2 Mutual Information Estimation
To estimate the mutual information between two random variables $U$ and $V$, we compute their Gram matrices $G_{U}$ and $G_{V}$, and form the joint similarity matrix via element-wise (Hadamard) product: $G_{UV}=G_{U}\circ G_{V}.$ After trace-normalization, mutual information is estimated by: I(U;V)=H(U)+H(V)H(U,V),I(U;V)=H(U)+H(V)-H(U,V), where $H(U,V)=-\text{tr}(G_{UV}\log G_{UV})$. The eigenvalue spectrum of $G_{UV}$ governs the joint entropy term; its shape reflects how much of the structure in $U$ and $V$ aligns. A more concentrated spectrum in $G_{UV}$ relative to $G_{U}$ and $G_{V}$ implies stronger dependence and thus higher mutual information.
A.3 Min Wasserstein Generalization Bound (He et al., 2024)
We restate the generalization bound proposed by He et al. (He et al., 2024), which characterizes generalization in terms of the Wasserstein distance between internal representations.
Suppose that the loss function $\tilde{\ell}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is $\rho_{0}$-Lipschitz, and the activation function $\phi_{\ell}:\mathbb{R}\to\mathbb{R}$ is $\rho_{\ell}$-Lipschitz for each $\ell=1,\ldots,L$. Then:
$\mathrm{gen}(P_{W|\mathcal{D}_{n}}\!,\!P_{X,Y})\!\leq\!\min_{\ell}\frac{\rho_{0}}{n}\sum_{i=1}^{n}\!\mathbb{E}_{W}\!\left[\!\left(1\!\vee\!\prod_{j=\ell+1}^{L}\!\rho_{j}\|W_{j}\|_{\mathrm{op}}\!\right)\!\cdot\!\mathcal{W}_{1}\!\left(\!P_{T_{\ell},i,Y_{i}|W}(\cdot|W),\!\ P_{T_{\ell},Y|W}(\cdot|W)\!\right)\!\right]$ (3)
This result shows that generalization can be tightly controlled by the Wasserstein distance (Villani et al., 2008) between representations at a specific layer—referred to the generalization funnel layer.
@@
The Min Wasserstein Generalization Bound (He et al., 2024) provides a theoretical foundation for our study by characterizing generalization in terms of distributional alignment at an intermediate layer. It motivates our analysis of information flow by suggesting that information is peaked at a specific layer—the generalization funnel. Our InfoRidge builds on this insight by quantifying predictive information across layers, and reveals that a specific intermediate layer exhibits peak mutual information and correlates with better generalization performance.
@@ B.2 CLUTRR
CLUTRR (Compositional Language Understanding and Text-based Relational Reasoning) (Sinha et al., 2019) is a diagnostic benchmark for evaluating relational inference in language models. Each example contains a story describing family relations, and the task is to infer the missing relationship between two entities. The distribution shift stems from clause lengths that are absent in the training set but present during evaluation. We use the task split “gen_train23_test2to10”, where the model is trained on clause lengths 2 and 3 and evaluated on lengths 2 through 10.
@@ B.3 ECQA
ECQA (Explanations for CommonsenseQA) (Aggarwal et al., 2021) is a commonsense multiple-choice question-answering dataset, where each question is accompanied by 5 answer options.
@@ B.4 Synthetic Arithmetic Dataset
$\texttt{S\{signal\}\_N\{noise\}},\quad\text{where noise}\sim\mathcal{U}_{\text{int}}(0,\texttt{noise\_range-1}).$
@@ C.1 Prompt Construction
All tasks are cast into a next-token generation format. The model receives a prompt and generate the next token. Below are construction strategies and examples for each dataset:
@@ E. Drive
C.2 Finetuning Settings
We fine-tuned all layers end-to-end using AdamW (learning rate $5\times 10^{-6}$, weight decay $0.01$) with a linear schedule (warmup ratio $0.1$) and early stopping on validation loss. Training converged in all settings.
C.3 Information Estimation
To estimate mutual information, we subsample between 50 and 200 test examples depending on model and task to achieve a stable result.
C.4 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
@@
  • $\beta$ are initialized to 1 and updated using gradient descent without altering the architecture.
@@
Table 4: Performance (%) of different models across datasets.
Model CLUTRR Synthetic ECQA
GPT-2 Small 30.98 71.45
GPT-2 Medium 34.21 68.35
Qwen-2.5-0.5B 31.94 83.20 57.06
LLaMA-3.1-8B 57.68 94.00 77.39
@@ Appendix D Compute and Licensing Details
Experiments were conducted on an NVIDIA RTX A6000 GPU (48GB).
@@
The GPT-2 Small (Radford et al., 2019) and GPT-2 Medium (Radford et al., 2019) models are released under the MIT License and are available via Hugging Face Transformers. The Qwen-2.5 0.5B (Yang et al., 2024) model is provided by Alibaba Group under the Apache 2.0 License. The LLaMA-3.1 8B (Meta AI, 2024) model is made available by Meta under a non-commercial research license and accessed via Hugging Face.
@@
CLUTRR (Sinha et al., 2019) is released under a CC BY 4.0 license as part of EMNLP 2019. The ECQA (Aggarwal et al., 2021) dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021. The Synthetic Arithmetic dataset is custom-designed by the authors and does not rely on any external or licensed data sources.
@@ Appendix E Full Quantitative Results with Confidence Estimates
To provide a comprehensive view of model behavior, we report the full experimental results for all models and datasets. This includes predictive information trends ($I(Z;Y)$), incremental information gain ($I(\Delta Z;Y)$) and residual scaling effects ($\beta_{\ell}$), with statistical significance reporting.
@@ E.1 Predictive Information
To estimate mutual information between hidden representations and target labels, we compute matrix-based mutual information following the methodology in Appendix A. We subsample between 50 and 100 test examples per experiment and report results with statistical significance based on multiple random seeds.
@@
Figure 7: Predictive information $I(Z;Y)$ across different models and datasets exhibits an information peak, indicating a generalization ridge. In cases where the task is too simple relative to model capacity—such as the synthetic arithmetic task with LLaMA—this trend reflects an overfitting regime. Lighter line colors represent later training epochs. Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.